OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation x + 2y = 4. x-intercept: (0, 2), y-intercept: (4, 0) x-intercept: (2, 0), y-intercept: (4, 0) x-intercept: (0, 4), y-intercept: (2, 0) x-intercept: (4, 0), y-intercept: (0, 2) Okay guys, I did this one in my head, and I want to double check it with some math genius'... I got the fourth solution... :) THANK YA'LL MUCH!

OpenStudy (amistre64):

x + 2y = 4 ; divde it all by 4 x/4 + 2y/4 = 1 ; simplify x/4 + y/2 = 1 when x=0; y = 2 when y = 0; x = 4 .. so your right

OpenStudy (anonymous):

the fourth one is correct.

OpenStudy (anonymous):

Thanks guys, I just wanted to be sure!

OpenStudy (amistre64):

\[\color{purple}{\text{youre welcome :)}}\]